OpenStudy (anonymous):

SoLvE the matrix equation: 2x-5[2 -8 -4 2] = [4 -6 2 -8]

OpenStudy (anonymous):

so the 2x-5 is getting multiplied into the first matrix? and then we solve for x?

jimthompson5910 (jim_thompson5910):

Is it this \[\Large 2X + 5\begin{bmatrix}2 & -8\\-4 & 2\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}\] ???

OpenStudy (anonymous):

yeah im trying to see if thats it..

OpenStudy (anonymous):

yes it is. exactly like tht lol

OpenStudy (anonymous):

or if it's (2x+5) time each of the elements as a scalar

OpenStudy (anonymous):

which would be painful..

jimthompson5910 (jim_thompson5910):

Start by multiplying the 5 through to each element in the matrix \[\Large 2X - 5\begin{bmatrix}2 & -8\\-4 & 2\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}\] \[\Large 2X - \begin{bmatrix}2*5 & -8*5\\-4*5 & 2*5\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}\] \[\Large 2X - \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}\] What's next?

OpenStudy (anonymous):

do i multiply everything by 2 next...

OpenStudy (anonymous):

sorry idk

jimthompson5910 (jim_thompson5910):

that will be one of your last steps, the next step is to add that matrix you see on the left side to both sides like this \[\Large 2X - \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}\] \[\Large 2X - \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} + \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}+ \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} \] \[\Large 2X + 0 = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}+ \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} \] \[\Large 2X = \begin{bmatrix}4 & -6\\2 & -8\end{bmatrix}+ \begin{bmatrix}10 & -40\\-20 & 10\end{bmatrix} \]

jimthompson5910 (jim_thompson5910):

Now add the two matrices on the right side

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

OpenStudy (anonymous):

[14 -46 -18 2]

jimthompson5910 (jim_thompson5910):

good, so \[\Large 2X = \begin{bmatrix}14& -46\\-18 & 2\end{bmatrix}\] the last step is to multiply both sides by 1/2 to completely isolate X \[\Large X = \frac{1}{2}\begin{bmatrix}14& -46\\-18 & 2\end{bmatrix}\] \[\Large X = \begin{bmatrix}\frac{1}{2}*14& \frac{1}{2}(-46)\\\frac{1}{2}(-18) & \frac{1}{2}*2\end{bmatrix}\] \[\Large X = \begin{bmatrix}7& -23\\-9 & 1\end{bmatrix}\]